# Length-Aware DETR for Robust Moment Retrieval

### v1

@@ 5 Limitation and Conclusion
Appendix A Additional Ablation Studies
Effect of MomentMix. We propose MomentMix, a novel mixing-based augmentation for moment retrieval (MR). To validate its effectiveness, we compared MomentMix with three naive augmentations: (1) Random Crop: randomly crops foreground regions. (2) Random Drop: masks a random $\alpha$of frame features by setting them to zero. (3) Gaussian Noise: adds Gaussian noise with mean=0 and std=$\beta$ to all frame features.
Table A1: Performance comparison of MomentMix and naive augmentations—Random Crop, Random Drop ($\alpha$, drop rates), and Gaussian Noise ($\beta$, noise levels)—on the QVHighlights val set with QD-DETR as the baseline. Naive augmentations fail to improve performance, while MomentMix achieves significant gains in short moments and overall robustness in moment retrieval tasks.
Method Short Middle Long All
R1 mAP R1 mAP R1 mAP R1 mAP
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Random Crop 5.31 8.99 39.34 43.03 41.01 46.00 40.95 41.00
Random Drop (0.1) 4.41 7.47 39.06 44.37 43.05 47.58 41.28 41.44
Random Drop (0.3) 3.52 7.99 38.16 41.74 43.62 46.93 40.69 40.14
Random Drop (0.5) 4.15 7.24 39.87 44.05 42.44 47.24 41.48 40.95
Gaussian Noise (0.01) 4.38 5.99 38.45 42.18 44.65 48.77 41.49 40.53
Gaussian Noise (0.05) 3.05 6.40 38.39 42.84 41.13 46.26 39.78 40.32
Gaussian Noise (0.1) 2.66 5.01 36.22 40.64 42.94 48.74 39.00 39.35
MomentMix 12.26 16.44 42.89 47.69 45.78 52.75 46.83 47.84
The results in Table A1 show that while Random Crop slightly enhances performance on short moments, it degrades mAP for other lengths, diminishing overall effectiveness. Similarly, Random Drop and Gaussian Noise fail to produce meaningful performance gains. In contrast, MomentMix achieved notable improvements for short moments while also enhancing performance across all lengths, leading to substantial gains in overall metrics. These results demonstrate the effectiveness of MomentMix augmentation compared with other naive augmentations.
Effect of Length-Aware Decoder. We present the Length-Aware Decoder (LAD), a novel framework designed to improve moment center predictions by conditioning on moment length. Unlike Group-DETR [6] in object dectection, which employs a group-wise one-to-many matching strategy without explicit group definitions, LAD utilizes length class-wise one-to-one matching to generate length-wise expert queries. To validate LAD’s effectiveness as a specialized and robust solution for MR, we applied both LAD and Group-DETR to a shared baseline (QD-DETR).
Method Short Middle Long All
R1 mAP R1 mAP R1 mAP R1 mAP
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Group-DETR 4.90 3.84 40.24 40.72 43.14 43.79 42.17 37.97
LAD 6.99 10.26 39.48 44.68 43.34 51.23 42.36 43.91
@@ Appendix A Additional Ablation Studies
Method Short Middle Long All
@@
Method Short Middle Long All
@@ Appendix B Moment Length Class Selection
Method Short Middle Long All
@@
$\mathcal{N}_{c}=4$ 6.99 10.26 39.48 44.68 43.34 51.23 42.36 43.91
@@
+2.26 +4.47 +5.62 +3.92 +0.53 +3.84
Table A7: Performance comparison on the QVHighlights val set using additional audio modality. UVCOM results are reproduced from the original repository.
Method Short Middle Long All
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
UVCOM 4.45 11.00 44.18 48.03 43.89 48.84 64.26 49.42 44.76 64.92 45.29 44.70
+8.93 +6.77 +1.33 +3.09 +1.95 +4.67 +2.45 +3.55 +4.01 +3.12 +6.23 +5.40
[…]
Method Short Middle Long All
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
UVCOM 5.64 10.83 48.96 51.12 49.16 51.71 70.13 54.97 49.99 67.95 47.88 47.56
+9.19 +8.66 +0.27 +3.48 +0.40 +3.96 +0.97 +3.03 +2.86 +3.50 +6.96 +5.69

### v2

@@ 5 Discussion
Appendix A Additional Ablation Studies
Representation of augmented data. We analyzed the performance degradation of short moments from a data perspective and observed that short-moment visual features exhibit limited diversity. To address this, we applied our MomentMix augmentation strategy to enhance feature diversity. To verify whether MomentMix effectively resolved this issue, we visualized the feature distributions of augmented short moments. As shown in Fig. A1, we found that the previously limited short-moment feature diversity increased significantly, approaching levels comparable to non-short moments. This confirms that MomentMix successfully mitigates the feature diversity issue, leading to improved model performance.
Figure A1: We performed the same analysis as in Fig. 2 on training data with our augmentation. The augmented short moment features lie within the same distribution as the original training features but exhibits greater diversity.
Effect of Length-Aware Decoder. We present the Length-Aware Decoder (LAD), a novel framework designed to improve moment center predictions by conditioning on moment length. Unlike Group-DETR [4] in object dectection, which employs a group-wise one-to-many matching strategy without explicit group definitions, LAD utilizes length class-wise one-to-one matching to generate length-wise expert queries, as shown in Tab. A1. To validate LAD’s effectiveness as a specialized and robust solution for MR, we applied both LAD and Group-DETR to a shared baseline [25].
Figure A2: [Left] Group-DETR [4] employs one-to-many matching, where the same labels are utilized across all groups. [Right] Our length-wise matching is one-to-one, and it operates within each length class. By matching only the predictions and ground truths that belong to the same class, this approach enables the creation of length-wise expert queries.
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Group-DETR 4.90 3.84 40.24 40.72 43.14 43.79 42.17 37.97
LAD 8.76 11.01 40.55 45.53 43.69 50.76 43.65 44.48
@@ Appendix A Additional Ablation Studies
Method Short Middle Long Full
@@
Method Short Middle Long Full
@@ Appendix B Moment Length Class Selection
Method Short Middle Long Full
@@
$\mathcal{N}_{c}=4$ 8.76 11.01 40.55 45.53 43.69 50.76 43.65 44.48
Table A5: Performance comparison on the QVHighlights val set using additional audio modality. ‡ means the result reproduced from the original repository.
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
$^{\textbf{{(+8.93)}}}$ $^{\textbf{{(+6.77)}}}$ $^{\textbf{{(+1.33)}}}$ $^{\textbf{{(+3.09)}}}$ $^{\textbf{{(+1.95)}}}$ $^{\textbf{{(+4.67)}}}$ $^{\textbf{{(+2.45)}}}$ $^{\textbf{{(+3.55)}}}$ $^{\textbf{{(+4.01)}}}$ $^{\textbf{{(+3.12)}}}$ $^{\textbf{{(+6.23)}}}$ $^{\textbf{{(+5.40)}}}$
Table A6: Performance comparison on the QVHighlights val set using the $\text{InternVideo2}_{s2}\text{-6B}$ features for both video and text modalities. ‡ means the result reproduced from the original repository.
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
$^{\textbf{{(+9.19)}}}$ $^{\textbf{{(+8.66)}}}$ $^{\textbf{{(+0.27)}}}$ $^{\textbf{{(+3.48)}}}$ $^{\textbf{{(+0.40)}}}$ $^{\textbf{{(+3.96)}}}$ $^{\textbf{{(+0.97)}}}$ $^{\textbf{{(+3.03)}}}$ $^{\textbf{{(+2.86)}}}$ $^{\textbf{{(+3.50)}}}$ $^{\textbf{{(+6.96)}}}$ $^{\textbf{{(+5.69)}}}$
@@
$^{\textbf{{(+2.26)}}}$ $^{\textbf{{(+4.47)}}}$ $^{\textbf{{(+5.62)}}}$ $^{\textbf{{(+3.92)}}}$ $^{\textbf{{(+0.53)}}}$ $^{\textbf{{(+3.84)}}}$